# Stateful Large Language Model Serving with Pensieve

### v2

@@ Abstract.
In this paper, we design $Pensieve$, a system optimized for multi-turn conversation LLM serving. $Pensieve$ maintains the conversation state across requests by caching previously processed history to avoid duplicate processing. $Pensieve$’s multi-tier caching strategy can utilize both GPU and CPU memory to efficiently store and retrieve cached data. $Pensieve$ also generalizes the recent PagedAttention kernel to support attention between multiple input tokens with a GPU cache spread over non-contiguous memory. Our evaluation shows that $Pensieve$ can achieve 13-58% more throughput compared to vLLM and TensorRT-LLM and significantly reduce latency.
@@ 1. Introduction
We have built Pensieve and compared its performance against vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2023), state-of-the-art serving systems that do not cache state across requests. Experiments show that Pensieve can improve serving throughput by 13% to 58% and also significantly reduce latency at moderate load.
@@ 3.2. Challenges
Figure 6. Multi-token attention with non-contiguous KV cache. The batch contains two requests, whose tokens (Query representations) are shown on the left. The right side shows the requests’ context (KV-tokens) which reside in non-contiguous GPU memory (block 3, 1 for request 0 and block 0, 4, 5, 2, 7 for request 1).
@@ 4.2. A unified batch scheduler
Pensieve’s iteration-level scheduler is “clocked” for action by the completion of a generation step. In particular, after each iteration of token generation, the worker returns any finished request that has either emitted the end-of-sentence token or reached maximum decoding length. The scheduler finds the next request in its wait queue to join the batch, if there is room (i.e. the number of total tokens in the batch is fewer than some pre-configured threshold). We use the simple first-come-first-serve scheduling policy when choosing which new requests to join the batch.
@@ 4.4. Multi-token attention for non-contiguous cache
Our new kernel is similar to that of vLLM to the extent that both kernels need to support loading KV cache from non-contiguous GPU global memory into on-chip shared memory. Their main difference is illustrated in Figure 9. As vLLM’s kernel performs attention between a request’s single input token and the existing KV-tokens, its underlying computation can be described as two matrix-vector multiplication operations, as shown in Figure 9  [^1] The batched version of PagedAttention performs batched matrix-vector multiplications.. In contrast, our kernel handles multiple input tokens for each request, computing attention scores between all pairs of input tokens and the conversation’s existing KV-tokens. Therefore, its underlying computation can be described as two matrix-matrix multiplication operations, and the batched version of our kernel performs batched matrix-matrix multiplications. Like vLLM, we fuse the two multiplication operations according to (Dao et al., 2022). Because of the additional dimension in the Q tensor, our kernel has more parallelization and tiling opportunities on the GPU. However, care must be taken to handle the new challenge that different requests in the batch have different numbers of input tokens.
@@ 5. Implementation
We have implemented our prototype serving system Pensieve with $\sim$7K lines of C++/CUDA code. Pensieve manages KV cache and auxiliary data structure needed by multi-token attention on the GPU, but relies on PyTorch (v0.2.0, CUDA 11.8) C++ front-end APIs to execute GPU operators in Large Language Models. Based on PyTorch’s fused memory efficient attention, we develop our own fused multi-token attention kernel using NVIDIA Cutlass library to support performing attention with KV-tokens that reside in non-contiguous memory.
@@ 6.1. Experimental Setup
We evaluate Pensieve on Azure NC A100 v4 series, which are equipped with up to 4 A100-80GB GPUs, a 24-core AMD EPYC 7003 processor per GPU, and 220 GB CPU memory. For each system evaluated, we configure it to use 40GB GPU memory for KV cache for a fair comparison.
@@
We evaluate two different sizes for each model: a small one on a single GPU, and a large one partitioned onto 4 GPUs using Tensor Parallelization as done in Megatron-LM (Shoeybi et al., 2019). Detailed model hyper-parameters can be found in Table 1. By default, Llama 2 only uses GQA for models with over 70 billion parameters. To demonstrate Pensieve’s effectiveness when used with GQA, we changed the number of KV heads of Llama 2-13B from 40 to 10. In all experiments, the 16-bit half-precision float format is used for both model parameters and intermediate hidden representations.
@@
Baselines
We evaluate Pensieve against two state-of-the-art serving systems: vLLM (Kwon et al., 2023) (v0.2.0) and TensorRT-LLM (NVIDIA, 2023) (v0.9.0). Both of them use a stateless serving API. For each request, the new user prompt is appended to the history and then processed as an input request. vLLM uses PyTorch as its execution backend. By contrast, TensorRT-LLM compiles and optimizes the model using graph rewriting optimizations such as operator fusion, and executes the optimized model using the TensorRT Runtime. Since vLLM is an execution engine without a serving loop that drives the engine to process incoming requests, we implement such a driver that adds newly arrived requests into vLLM job queue and invokes the engine execution until all requests are fully processed.
@@ 6.2. End-to-End Serving Performance
Figure 10 shows the normalized latency vs throughput for OPT-13B and Llama 2-13B served with a single A100-80GB GPU. The normalized latency is calculated as the mean of each request’s end-to-end latency divided by its output length, as done in (Yu et al., 2022; Kwon et al., 2023). On the ShareGPT dataset shown in Figure 10(a) and (b), Pensieve achieves 33% more throughput over vLLM and 19% more over TensorRT-LLM for serving OPT-13B, and  57% more throughput over both vLLM and TensorRT-LLM for serving Llama 2-13B. On the UltraChat dataset shown in Figure 10(c) and (d), Pensieve achieves 17% and 13% more throughput over vLLM and TensorRT-LLM respectively for serving OPT-13B, 58% and 47% more throughput for serving Llama 2-13B. Pensieve has more throughput gains on ShareGPT than UltraChat because the real world dataset ShareGPT has more conversation turns (Table 2), which makes saving past KV-tokens more beneficial.
Figure 10 shows that TensorRT-LLM outperforms vLLM. This is expected because TensorRT-LLM compiles and optimizes the model before execution. However, despite the fact that Pensieve also executes the model directly using Torch like vLLM, it outperforms TensorRT-LLM by avoiding recomputing past KV-tokens for continuing conversations.
Pensieve’s performance gain is more significant for Llama 2-13B than for OPT-13B because our version of Llama 2-13B uses Grouped-Query Attention with group size 4 (i.e. every four query heads share a single key-value head). Consequently, the amount of memory required to store past KV-tokens is reduced by 4x, thus allowing Pensieve to store more past KV-tokens and better avoid recomputation.
@@ 6.3. Multi-GPU Serving Performance
Figure 11 shows the performance of Pensieve for larger models, OPT-66B and Llama 2-70B, when run on four GPUs using the ShareGPT dataset. Larger models amplify Pensieve’s advantage over the baselines because the amount of computation grows faster than the memory usage of KV-tokens. For example, from OPT-13B to OPT-66B, the model parameter size and computation amount increase by more than 4x, while the hidden size only increases 1.8x from 5120 to 9216 (Table 1). Since the number of GPUs and CPU memory are usually scaled linearly with the model size, Pensieve can store more past KV-tokens in its KV cache.
Pensieve achieves 1.5x the throughput of baseline systems for OPT-66B and 1.6x for Llama 2-70B. The improvement is more significant on Llama 2-70B because it uses Grouped-Query Attention with group size 8, which reduces the memory requirement for past KV-tokens by 8x. This much-reduced memory consumption in KV-tokens also significantly benefits the throughput of Pensieve (GPU cache), as shown in Figure 11(b).
@@ 6.4. Performance of Multi-token attention Kernel
Figure 12. Execution time of Pensieve’s multi-token attention kernel over non-contiguous context memory with different context size. Measured with batch size 32 and query size 8.
@@ 6.6. Effect of the Eviction Policy
We compare Pensieve’s caching policy against the classic LRU policy. We use OPT-13B as the workload for this evaluation. From Figure 14 shows that both policies exhibit similar performance until the workload approaches 3 requests per second, beyond which Pensieve’s eviction policy outperforms LRU. After analyzing the execution traces, we find that both policies have less than 80% cache hit rate, however, Pensieve’s policy has up to 4.4 percentage points higher CPU cache hit rate than LRU. On average, Pensieve’s policy reduces the number of recomputed KV-tokens by up to 14.6%.
@@ 7. Related Works
The system of  (Pope et al., 2023) can keep a fixed-sized KV cache for each conversation in TPU device memory. Such caching is too rigid. By contrast, Pensieve manages the cache across all conversations and can also swap to CPU to relieve GPU memory pressure. Another recent system, CacheGen (Liu et al., 2023), addresses the problem of reducing the context-loading delay for long contexts stored in remote network storage. CacheGen proposes an adaptive compression scheme to accelerate the transfer of the KV cache over a bandwidth-limited and bursty network. As Pensieve only keeps the KV cache in local GPU and CPU memory, CacheGen’s techniques are orthogonal to those of Pensieve. For conversational chatbots, Pensieve using local cache can already bring significant performance gains.
@@ 8. Conclusion
When serving Large Language Models for multi-turn conversations, a major inefficiency is due to the recomputation of a conversation’s cumulative history context. We develop Pensieve, a stateful LLM serving system that preserves history embeddings in a multi-tier GPU-CPU cache. It uses a new GPU attention kernel to perform attention between requests’ new multi-token input and their saved context stored in non-contiguous GPU memory. Experiments show that Pensieve improves serving throughput by 13-58% compared to the state-of-the-art systems vLLM and TensorRT-LLM.

### v3

@@ Abstract.
In this paper, we design $Pensieve$, a system optimized for multi-turn conversation LLM serving. $Pensieve$ maintains the conversation state across requests by caching previously processed history to avoid duplicate processing. $Pensieve$’s multi-tier caching strategy can utilize both GPU and CPU memory to efficiently store and retrieve cached data. $Pensieve$ also generalizes the recent PagedAttention kernel to support attention between multiple input tokens with a GPU cache spread over non-contiguous memory. Our evaluation shows that $Pensieve$ can achieve $1.14$-3.0×3.0$\times$3.0 × the throughput of vLLM and TensorRT-LLM and significantly reduce latency.
LLM Serving, Multi-turn Conversations, Cache
journalyear: 2025copyright: acmlicensedconference: Twentieth European Conference on Computer Systems; March 30–April 3, 2025; Rotterdam, Netherlandsbooktitle: Twentieth European Conference on Computer Systems (EuroSys ’25), March 30–April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3689031.3696086isbn: 979-8-4007-1196-1/25/03submissionid: 627ccs: Software and its engineering Virtual memoryccs: Computer systems organization Data flow architectures
@@ 1. Introduction
We have built Pensieve and compared its performance against vLLM (Kwon et al., 2023) and TensorRT-LLM (NVIDIA, 2023), two state-of-the-art serving systems that do not cache state across requests. Experiments show that Pensieve achieves serving throughput that is $1.14$-1.70×1.70$\times$1.70 × of vLLM and TensorRT-LLM for small single-GPU models (OPT-13B, Llama 2-13B) and $1.64$-3.0×3.0$\times$3.0 × for large four-GPU models (OPT-66B, Llama 2-70B). At moderate serving load, Pensieve also significantly reduces latency.
@@ 3.2. Challenges
Figure 6. Multi-token attention with non-contiguous KV cache. The batch contains two requests, whose tokens (Query representations) are shown on the left. The right side shows the requests’ context (KV-tokens) which reside in non-contiguous GPU memory (chunk 3, 1 for request 0 and chunk 0, 4, 5, 2, 7 for request 1).
@@ 4.2. A unified batch scheduler
Pensieve’s iteration-level scheduler is “clocked” for action by the completion of a generation step. In particular, after each iteration of token generation, the worker returns any finished request that has either emitted the end-of-sentence token or reached maximum decoding length. The scheduler finds the next request in its wait queue to join the batch, if there is room (i.e. the total number of tokens in the batch is fewer than some pre-configured threshold). We use the simple first-come-first-serve scheduling policy when choosing which new requests to join the batch.
@@ 4.4. Multi-token attention for non-contiguous cache
Our new kernel is similar to that of vLLM to the extent that both kernels need to support loading KV cache from non-contiguous GPU global memory into on-chip shared memory. Their main difference is illustrated in Figure 9. As vLLM’s kernel performs attention between a request’s single input token and the existing KV-tokens, its underlying computation can be described as two matrix-vector multiplication operations [^1] The batched version of PagedAttention performs batched matrix-vector multiplications., as shown in Figure 9. In contrast, our kernel handles multiple input tokens for each request, computing attention scores between all pairs of input tokens and the conversation’s existing KV-tokens. Therefore, its underlying computation can be described as two matrix-matrix multiplication operations, and the batched version of our kernel performs batched matrix-matrix multiplications. Like vLLM, we fuse the two multiplication operations according to (Dao et al., 2022). Because of the additional dimension in the Q tensor, our kernel has more parallelization and tiling opportunities on the GPU. However, care must be taken to handle the new challenge that Q is a ragged-sized tensor as different requests in the batch have different numbers of input tokens.
@@ 4.4.1. Unifying the prefill and generation phase
4.4.2. Multi-GPU support
For large LLM models that do not fit into a single GPU, we adopted Tensor Parallelization, as implemented in Megatron-LM (Shoeybi et al., 2019), to partition large models across multiple GPUs. The KV cache in Pensieve is also partitioned accordingly. Each model partition is managed by a worker process which has its own GPU and CPU cache partition to store the attention state partition assigned to this worker. Since the model partitioning occurs along the feature dimension of KV cache, it does not affect the eviction policy that decides which token to migrate or drop. Therefore, each worker follows the same migration plan to swap attention states between GPU and CPU memory.
@@ 5. Implementation
We have implemented our prototype serving system Pensieve with $\sim$7K lines of C++/CUDA code. Pensieve manages KV cache and auxiliary data structure needed by multi-token attention on the GPU, but relies on PyTorch (v2.0.0, CUDA 11.8) C++ front-end APIs to execute GPU operators in Large Language Models. Based on PyTorch’s fused memory efficient attention, we develop our own fused multi-token attention kernel using NVIDIA Cutlass library to support performing attention with KV-tokens that reside in non-contiguous memory.
@@ 6.1. Experimental Setup
We evaluate Pensieve on Azure NC A100 v4 series, which are equipped with up to 4 A100-80GB GPUs, a 24-core AMD EPYC 7003 processor and 220 GB CPU memory per GPU. For each system evaluated, we configure it to allocate 40 GB memory on each GPU for KV cache for a fair comparison.
@@
We evaluate two different sizes for each model: a small one on a single GPU, and a large one partitioned onto 4 GPUs using Tensor Parallelization as done in Megatron-LM (§4.4.2). Detailed model hyper-parameters can be found in Table 1. By default, Llama 2 only uses GQA for models with over 70 billion parameters. To demonstrate Pensieve’s effectiveness when used with GQA, we changed the number of KV heads of Llama 2-13B from 40 to 10. In all experiments, the 16-bit half-precision float format is used for both model parameters and intermediate hidden representations.
@@
Baselines
We evaluate Pensieve against two state-of-the-art serving systems: vLLM (Kwon et al., 2023) (v0.2.0) and TensorRT-LLM (NVIDIA, 2023) (v0.12.0). Both of them use a stateless serving API. For each request, the new user prompt is appended to the history and then processed as an input request. vLLM uses PyTorch as its execution backend. By contrast, TensorRT-LLM compiles and optimizes the model using graph rewriting optimizations such as operator fusion, and executes the optimized model using the TensorRT Runtime. Since vLLM is an execution engine without a serving loop that drives the engine to process incoming requests, we implement such a driver that adds newly arrived requests into vLLM job queue and invokes the engine execution until all requests are fully processed.
@@ 6.2. End-to-End Serving Performance
Figure 10 shows the normalized latency vs throughput for OPT-13B and Llama 2-13B served with a single A100-80GB GPU. The normalized latency is calculated as the mean of each request’s end-to-end latency divided by its output length, as done in (Yu et al., 2022; Kwon et al., 2023). On the ShareGPT dataset shown in Figure 10(a) and (b), Pensieve’s throughput is 1.36×1.36$\times$1.36 × of vLLM and 1.14×1.14$\times$1.14 × of TensorRT-LLM for serving OPT-13B at 120 ms per token latency, and 1.70×1.70$\times$1.70 × of vLLM and 1.58×1.58$\times$1.58 × of TensorRT-LLM for serving Llama 2-13B at 180 ms per token latency. On the UltraChat dataset shown in Figure 10(c) and (d), Pensieve’s throughput is 1.17×1.17$\times$1.17 × and 1.14×1.14$\times$1.14 × of vLLM and TensorRT-LLM respectively for serving OPT-13B at 120 ms per token latency, 1.58×1.58$\times$1.58 × and 1.47×1.47$\times$1.47 × for serving Llama 2-13B at 100 ms per token latency. Pensieve has more throughput gains on ShareGPT than UltraChat because the real world dataset ShareGPT has more conversation turns (Table 2), which makes saving past KV-tokens more beneficial.
In Figure 10, TensorRT-LLM outperforms vLLM consistently. It even outperforms Pensieve (GPU cache) in Figure 10(a) and (c). This is because both Pensieve and vLLM use PyTorch API to execute the model while TensorRT-LLM performs optimizations like graph rewriting and operator fusion during an offline compiling stage, and then executes the optimized model using the TensorRT Runtime. However, by avoiding recomputing past KV-tokens for continuing conversations, Pensieve achieves significant throughput gain over TensorRT-LLM.
Pensieve’s performance gain is more significant for Llama 2-13B than for OPT-13B because our version of Llama 2-13B uses Grouped-Query Attention with group size 4 (i.e. every four query heads share a single key-value head). Consequently, the amount of memory required to store past KV-tokens is reduced by 4×4$\times$4 ×, thus allowing Pensieve to store more past KV-tokens and better avoid recomputation.
@@ 6.3. Multi-GPU Serving Performance
Figure 11 shows the performance of Pensieve for larger models, OPT-66B and Llama 2-70B, when run on four GPUs using the ShareGPT dataset. Larger models amplify Pensieve’s advantage over the baselines because the amount of computation grows faster than the memory usage of KV-tokens. For example, from OPT-13B to OPT-66B, the model parameter size and computation amount increase by more than 5×5$\times$5 ×, while the KV cache size (# layer $\times$ # hidden $\times$ 2) only increases by 2.88×2.$8$ (Table 1). Since the number of GPUs and CPU memory are usually scaled linearly with the model size, Pensieve can store more past KV-tokens in its KV cache.
Pensieve achieves 2.04×2.04$\times$2.04 × the throughput of vLLM and 1.64×1.64$\times$1.64 × of TensorRT-LLM for serving OPT-66B at 200 ms per token latency, and 3.0×3.0$\times$3.0 × of vLLM and 2.47×2.47$\times$2.47 × of TensorRT-LLM for serving Llama 2-70B at 400 ms per token latency. The improvement is more significant on Llama 2-70B because it uses Grouped-Query Attention with group size 8, which reduces the memory requirement for past KV-tokens by 8×8$\times$8 ×. This much-reduced memory consumption in KV-tokens also significantly benefits the throughput of Pensieve (GPU cache), as shown in Figure 11(b).
@@ 6.4. Performance of Multi-token attention Kernel
Figure 12. Execution time of Pensieve’s multi-token attention kernel over non-contiguous context memory with different context sizes. Measured with batch size 32 and query size 8.
@@ 6.6. Effect of the Eviction Policy
We compare Pensieve’s caching policy against the classic LRU policy. We use OPT-13B as the workload for this evaluation. Figure 14 shows that both policies exhibit similar performance until the workload approaches 3 requests per second, beyond which Pensieve’s eviction policy outperforms LRU. After analyzing the execution traces, we find that both policies have less than 80% cache hit rate, however, Pensieve’s policy has up to 4.4 percentage points higher CPU cache hit rate than LRU. On average, Pensieve’s policy reduces the number of recomputed KV-tokens by up to 14.6%.
@@ 7. Related Works
PromptCache (Gim et al., 2024) SGLang (Zheng et al., 2023) CachedAttention (Gao et al., 2024) RAGCache (Jin et al., 2024) CacheGen (Liu et al., 2024) Pensieve
Cache shared across users Yes Yes No Yes Yes No
Cache storage CPU GPU CPU & Disk CPU & GPU Network CPU & GPU
Eviction granularity N/A requests entire conversation history documents N/A tokens
Eviction location preference N/A trailing N/A trailing N/A leading
Table 3. Comparison of Pensieve with other caching systems for LLM attention states.
Systems that cache LLM attention states.
Several systems (Table 3), developed concurrently with Pensieve, also cache attention states in LLM serving to reduce computation.
PromptCache (Gim et al., 2024), SGLang (Zheng et al., 2023), RAGCache (Jin et al., 2024) and CachedAttention (Gao et al., 2024) all cache attention states of shared text segments across different requests. Among these systems, CachedAttention targets multi-turn conversations, RAGCache targets retrieval-augmented generation (RAG) while others aim to be general. PromptCache requires users to define a schema a priori to explicitly specify any reuse opportunities. This requirement is too restrictive since static schemas cannot specify those dynamic prefixes to be cached in multi-turn conversations. SGLang and RAGCache employ a tree-based cache in order to enable prefix sharing across different users or different RAG document prefixes. By contrast, Pensieve and CachedAttention are designed specifically for multi-turn conversation scenarios where different users are unlikely to share conversational content [^3] If the chatbot uses a common system prompt across users, it can be handled by explicitly designating the system prompt state as reusable.. In this setting, there is no need for the extra complexity of a tree-based cache structure. Instead, Pensieve and CachedAttention both focus on managing a dynamic cache for ongoing conservation histories which grow in size over time.
As Table 3 shows, the important difference between Pensieve and others (SGLang (Zheng et al., 2023), RAGCache (Jin et al., 2024) and CachedAttention (Gao et al., 2024)) lies in Pensieve’s eviction strategy which evicts at the finer-granularity of groups of tokens at the leading end to reduce the cost of recomputation. By contrast, CachedAttention evicts at the granularity of an entire conversation and does not recompute truncated conversation context, while SGLang and RAGCache evict at the granularity of a tree node containing an entire user request or document. SGLang and RAGCache also preferentially evict from the bottom of the tree (aka the trailing end of a prefix) as opposed to the leading end as done in Pensieve. Pensieve’s eviction strategy introduces challenges in supporting computation in non-sequential regions of request context (Figure 5), which is addressed in Pensieve’s multi-token attention kernel (§4.3.4).
Another recent system, CacheGen (Liu et al., 2024), addresses the problem of reducing the context-loading delay for long contexts stored in remote network storage. CacheGen proposes an adaptive compression scheme to accelerate the transfer of the KV cache over a bandwidth-limited and bursty network. As Pensieve only keeps the KV cache in local GPU and CPU memory, CacheGen’s techniques are orthogonal to those of Pensieve. For conversational chatbots, Pensieve using local cache can already bring significant performance gains.
@@ 8. Conclusion
When serving Large Language Models for multi-turn conversations, a major inefficiency is due to the recomputation of a conversation’s cumulative history context. We develop Pensieve, a stateful LLM serving system that preserves history embeddings in a multi-tier GPU-CPU cache. It uses a new GPU attention kernel to perform attention between requests’ new multi-token input and their saved context stored in non-contiguous GPU memory. Experiments show that Pensieve achieves $1.14$-1.70×1.70$\times$1.70 × the throughput of baseline systems for small models (OPT-13B, Llama 2-13B) and $1.64$-3.0×3.0$\times$3.0 × for larger models (OPT-66B, Llama 2-70B).
Acknowledgements.
We would like to thank our shepherd Yizhou Shan and the anonymous EuroSys reviewers for their constructive feedback. This project is partially supported by the National Science Foundation under Grant No. 2220407 and through the NYU IT High Performance Computing resources, services, and staff expertise. Lingfan Yu is supported by an AMD research award.